# Diffusion Models Meet Remote Sensing: Principles, Methods, and Perspectives

### v1

@@ I Introduction
In summary, diffusion models possess great development potential in the field of RS. Therefore, it is necessary to review and summarize existing diffusion model-based RS papers to help researchers gain a comprehensive understanding of current research status, and identify the gaps in the application of diffusion models in RS, thereby promoting further development in this field.
@@ II Theoretical Background of Diffusion Models
Diffusion models, also known as diffusion probabilistic models, are a family of deep generative models. In general, generative model is to convert a random distribution (i.e., noise) into a “probability distributions” that is the same as the distribution of the observed dataset, and obtain the desired outcomes by sampling from this “probability distribution”.
Obviously, it is quite difficult to obtain the target distribution directly, but disrupting a regular distribution into random noise is straightforward, and can be achieved by continuously adding Gaussian noise, as illustrated in Fig. 2. Diffusion model is inspired by this process and realizes the generation of target data by learning the reverse denoising process of this process, which can be traced back to 2015 [42] and became popular after denoising diffusion probabilistic model (DDPM) was published in 2020 [15].
@@ III-A RS Image Generation
As one of the most impressive deep generative models, diffusion models are expected to synthesize realistic RS images from existing images or given textual descriptions to support the development of various RS applications. According to the data sources, these image generation methods can be mainly divided into two categories: text-to-image generation and image-to-image generation.
@@ III-A1 Text-to-Image
Over the past two years, numerous text-to-image diffusion models have come out in computer vision [49, 47, 48, 16], especially the Stable Diffusion (SD) [48] model that has been widely adopted since its release [50, 51, 52]. However, its success mainly depends on training with billions of text-image pairs from the internet [53], which makes it difficult to extend to the field of RS since such vast and diverse RS datasets are not readily available. To address this problem, a straightforward idea is to produce trainable RS text-image pairs. Ou et al. [54] realized this idea with the help of pre-trained large models. Specifically, they first caption the existing RS images through a vision-language pre-training model to obtain initial textual prompts. Then, refining these prompts with human feedback and GPT-4 to improve semantic accuracy and suitability, successfully enabling the SD model to synthesize the required RS images. Instead of directly supplementing text prompts, Khanna et al. [55] used various numerical information related to satellite images, such as geolocation and sampling time as new prompts of the SD model, which effectively enriches the SD model’s input and enhances its ability to generate high-quality satellite images. Furthermore, Tang et al. [56] refined the generation process with SD model by incorporating RS image-related features as control conditions. They treated textual descriptions and numerical information as global control information, and used the depth map, segmentation mask, object boundaries, and other result images obtained through a series of pre-trained networks as local control information. By flexibly selecting control conditions, this approach achieves effective integration of multiple control information, expanding the RS image generation space.
@@ III-B2 Cloud Removal
The control condition for cloudy image reconstruction comes in various forms. For example, Czerkawski et al. [101] adopted text prompts and edge information [102] as the guiding conditions, together with the input cloudy image, cloud mask, and diffused cloud-free image to control the generation process of SD [48]. Jing et al. [103] input both SAR image and cloudy optical RS image into the diffusion model for feature extraction, effectively enhancing the cloud removal results with the help of cloud-unaffected SAR image. Zou et al. [104] first extracted features from the cloudy image and noise level, and then inputs the extracted spatial and temporal features into the diffusion model as control conditions. Instead of using more control conditions, they trained the model in an supervised manner with cloud-free images to further improve the quality of the reconstructed images. Moreover, Zhao et al. [105] integrated different images from multi-modalities and multi-time into a sequence input, and utilized a two-branch diffusion model to extract scene content from the optical RS and SAR images, respectively. Unlike the previous works, this method does not require the participation of cloud-free images and can handle image sequences of any length, offering greater flexibility and practical value.
@@ III-B3 Denoising
Due to the inherent constraints of imaging technology and environmental conditions, RS images are always accompanied by various noise [106]. In essence, the learning process of diffusion model is equal to the denoising process [15], which makes it possible to achieve superior denoising performance in the context of RS. Different modalities of RS images confront different noise challenges. For example, optical RS images are more susceptible to blurring caused by atmospheric scattering and absorption [107, 108], as well as the noise from changing lighting conditions [109]. HSIs need to consider the uneven distribution of noise over the spectral dimensions [110] and the correlation of noise between different bands [111, 112]. Fortunately, these noises can be effectively removed through diffusion models. Huang et al. [107] proposed to crop the noisy RS image into small regions and rearrange them in a cyclic shift manner before fed into the diffusion model, so as to achieve finer local denoising as well as artifact elimination. He et al. [111] proposed a truncated diffusion model that starts denoising from the intermediate step of the diffusion process, instead of a pure noise, to avoid the destruction of the inherent effective information in the HSI. Moreover, Yu et al. [113] simulated the harsh imaging conditions of RS satellites by adding various attack disturbances to input images, enhancing the diffusion model’s ability to counteract the system noise.
@@ III-C1 Landcover Classification
As one of the most common applications in the RS community, landcover classification aims to assign each pixel to a specific class, such as buildings or grass, to obtain useful landcover information. However, the RS images always contain diverse and complex scenes, increasing the difficulty of accurate classification. Given that diffusion model can learn and simulate complex data distributions better than other deep learning models, researchers are trying to use it for the RS landcover classification.
@@ III-C3 Climate Prediction
Climate prediction is another application for diffusion model in the field of RS, which is a complex systematic project that requires the integration of multiple variables, including cloud amount, cyclone distribution, and water vapor. Therefore, its solution often consists of multiple diffusion models, enabling fine-grained, stepwise processing of these high-dimensional data. For example, Nath et al. [135] cascaded three independently trained diffusion models to generate future satellite imagery, increase the resolution, and predict precipitation. Hatanaka et al. [136] utilized two cascaded score-based diffusion models to generate high-resolution cloud cover images from coarse-resolution atmospheric variables. In addition, Leinonen et al. [137] addressed the issue of high computational costs by adopting the latent diffusion model concept proposed in [48], running the diffusion process within a latent variable space mapped by a 3D VAE network. These methods all demonstrate that diffusion models can better capture the complex spatio-temporal relationships and become a powerful tool in climate prediction.
@@ III-C4 Miscellaneous Tasks
One of such tasks is height estimation, which aims at providing pixel-wise height information of surface features (e.g., buildings, trees, terrain, etc.) to generate 3D models of surface scenes. Diffusion model has been reported to be a promising solution for surface estimation [138]. Unlike traditional methods that require multi-view geospatial imagery or LiDAR point clouds, it produces accurate height estimates only with single-view optical RS images.

### v2

@@ I Introduction
In summary, diffusion models possess great development potential in the field of RS. Therefore, it is necessary to review and summarize existing diffusion model-based RS papers to help researchers gain a comprehensive understanding of the current research status, and identify the gaps in the application of diffusion models in RS, thereby promoting further development in this field.
@@ II Theoretical Background of Diffusion Models
Diffusion models, also known as diffusion probabilistic models, constitute a family of deep generative models. In essence, a generative model aims to transform a random distribution (i.e., noise) into a ’probability distribution’ that matches the distribution of the observed dataset, thereby producing desired outcomes through sampling from this ’probability distribution’.
Obtaining the target distribution directly is evidently challenging. However, disrupting a regular distribution into random noise is comparatively straightforward and can be accomplished by continually adding Gaussian noise, as depicted in Fig. 2. The concept of diffusion models is inspired by this process and involves learning the reverse denoising process, which dates back to 2015 [42] and gained popularity following the publication of the denoising diffusion probabilistic model (DDPM) in 2020 [15].
@@ III-A RS Image Generation
As one of the most impressive deep generative models, diffusion models are expected to synthesize realistic RS images from existing images or given textual descriptions to support the development of various RS applications, as an alternative to autoregressive models [49]. According to the data sources, these image-generation methods can be mainly divided into two categories: text-to-image generation and image-to-image generation.
@@ III-A1 Text-to-Image
Over the past two years, numerous text-to-image diffusion models have come out in computer vision [47, 48, 16], especially the Stable Diffusion (SD) [48] model that has been widely adopted since its release [50, 51, 52]. However, its success mainly depends on training with billions of text-image pairs from the internet [53], which makes it difficult to extend to the field of RS since such vast and diverse RS datasets are not readily available, as mentioned in [49]. To address this problem, a straightforward idea is to produce trainable RS text-image pairs. Ou et al. [54] realized this idea with the help of pre-trained large models. Specifically, they first caption the existing RS images through a vision-language pre-training model to obtain initial textual prompts. Then, refining these prompts with human feedback and GPT-4 to improve semantic accuracy and suitability, successfully enabling the SD model to synthesize the required RS images. Instead of directly supplementing text prompts, Khanna et al. [55] used various numerical information related to satellite images, such as geolocation and sampling time as new prompts of the SD model, which effectively enriches the SD model’s input and enhances its ability to generate high-quality satellite images. Furthermore, Tang et al. [56] refined the generation process with SD model by incorporating RS image-related features as control conditions. They treated textual descriptions and numerical information as global control information, and used the depth map, segmentation mask, object boundaries, and other result images obtained through a series of pre-trained networks as local control information. By flexibly selecting control conditions, this approach achieves effective integration of multiple control information, expanding the RS image generation space.
@@ III-B2 Cloud Removal
The control condition for cloudy image reconstruction comes in various forms. For example, Czerkawski et al. [101] adopted text prompts and edge information [102] as the guiding conditions, together with the input cloudy image, cloud mask, and diffused cloud-free image to control the generation process of SD [48]. Jing et al. [103] input both SAR image and cloudy optical RS image into the diffusion model for feature extraction, effectively enhancing the cloud removal results with the help of cloud-unaffected SAR image. Zou et al. [104] first extracted features from the cloudy image and noise level, and then inputs the extracted spatial and temporal features into the diffusion model as control conditions. Instead of using more controlled conditions, they trained the model in a supervised manner with cloud-free images to further improve the quality of the reconstructed images. Moreover, Zhao et al. [105] integrated different images from multi-modalities and multi-time into a sequence input, and utilized a two-branch diffusion model to extract scene content from the optical RS and SAR images, respectively. Unlike the previous works, this method does not require the participation of cloud-free images and can handle image sequences of any length, offering greater flexibility and practical value.
@@ III-B3 Denoising
Due to the inherent constraints of imaging technology and environmental conditions, RS images are always accompanied by various noise [106]. In essence, the learning process of the diffusion model is equal to the denoising process [15], which makes it possible to achieve superior denoising performance in the context of RS. Different modalities of RS images confront different noise challenges. For example, optical RS images are more susceptible to blurring caused by atmospheric scattering and absorption [107, 108], as well as the noise from changing lighting conditions [109]. HSIs need to consider the uneven distribution of noise over the spectral dimensions [110] and the correlation of noise between different bands [111, 112]. Fortunately, these noises can be effectively removed through diffusion models. Huang et al. [107] proposed to crop the noisy RS image into small regions and rearrange them in a cyclic shift manner before fed into the diffusion model, so as to achieve finer local denoising as well as artifact elimination. He et al. [111] proposed a truncated diffusion model that starts denoising from the intermediate step of the diffusion process, instead of a pure noise, to avoid the destruction of the inherent effective information in the HSI. Moreover, Yu et al. [113] simulated the harsh imaging conditions of RS satellites by adding various attack disturbances to input images, enhancing the diffusion model’s ability to counteract the system noise.
@@ III-C1 Landcover Classification
As one of the most common applications in the RS community, landcover classification aims to assign each pixel to a specific class, such as buildings or grass, to obtain useful landcover information. However, the RS images always contain diverse and complex scenes, increasing the difficulty of accurate classification. Given that the diffusion model can learn and simulate complex data distributions better than other deep learning models, researchers are trying to use it for the RS landcover classification.
@@ III-C3 Climate Prediction
Climate prediction is another application for the diffusion model in the field of RS, which is a complex systematic project that requires the integration of multiple variables, including cloud amount, cyclone distribution, and water vapor. Therefore, its solution often consists of multiple diffusion models, enabling fine-grained, stepwise processing of these high-dimensional data. For example, Nath et al. [135] cascaded three independently trained diffusion models to generate future satellite imagery, increase the resolution, and predict precipitation. Hatanaka et al. [136] utilized two cascaded score-based diffusion models to generate high-resolution cloud cover images from coarse-resolution atmospheric variables. In addition, Leinonen et al. [137] addressed the issue of high computational costs by adopting the latent diffusion model concept proposed in [48], running the diffusion process within a latent variable space mapped by a 3D VAE network. These methods all demonstrate that diffusion models can better capture the complex spatio-temporal relationships and become a powerful tool in climate prediction.
@@ III-C4 Miscellaneous Tasks
One such task is height estimation, which aims at providing pixel-wise height information of surface features (e.g., buildings, trees, terrain, etc.) to generate 3D models of surface scenes. Diffusion model has been reported to be a promising solution for surface estimation [138]. Unlike traditional methods that require multi-view geospatial imagery or LiDAR point clouds, it produces accurate height estimates only with single-view optical RS images.